# Faces of Fairness: Examining Bias in Facial Expression Recognition Datasets and Models

### v1

@@ Abstract
Building AI systems, including Facial Expression Recognition (FER), involves two critical aspects: data and model design. Both components significantly influence bias and fairness in FER tasks. Issues related to bias and fairness in FER datasets and models remain underexplored. This study investigates bias sources in FER datasets and models. Four common FER datasets—AffectNet, ExpW, Fer2013, and RAF-DB—are analyzed. The findings demonstrate that AffectNet and ExpW exhibit high generalizability despite data imbalances. Additionally, this research evaluates the bias and fairness of six deep models, including three state-of-the-art convolutional neural network (CNN) models: MobileNet, ResNet, XceptionNet, as well as three transformer-based models: ViT, CLIP, and GPT-4o-mini. Experimental results reveal that while GPT-4o-mini and ViT achieve the highest accuracy scores, they also display the highest levels of bias. These findings underscore the urgent need for developing new methodologies to mitigate bias and ensure fairness in datasets and models, particularly in affective computing applications. See our implementation details at https://github.com/MMHosseini/bias_in_FER.
@@
Bias, Fairness, Facial Expression Recognition, Facial Affect Analysis, Bias in FER Datasets, Bias in FER Algorithms.
@@ 1 Introduction
Despite advancements in deep learning, bias and fairness issues in FER systems remain unexplored. Bias refers to prejudices or favoritism toward groups, while fairness ensures impartiality and avoids disparities [9, 10]. Bias can be analyzed in two dimensions: datasets and models [9, 11, 12]. Datasets mainly suffer from missing data, noisy data, labeling problems, imbalanced distribution, scalability, and bias. However, some classification approaches have been proposed to tackle these problems [13, 14, 15]. Dataset bias often arises from demographic imbalances, such as disparities in age, gender, or race [9]. Model bias stems from architecture, training process, and unrepresentative datasets [16, 17, 18, 19]. Bias impacts model performance for underrepresented groups, reduces generalizability, and compromises fairness. Addressing bias in datasets and models is crucial for fairness in FER.
Bias in machine learning can be classified in various ways. Pagano et al. [20] identified three types: (1) pre-existing bias from inequalities in data, (2) technical bias introduced during model design, and (3) emergent bias causing system-level disparities. Pagano’s model argues bias mainly originates from data [21]: underrepresented demographic groups in training data lead to poorer performance. Expanding this theory, Cheong et al. [22] identified three sources: i) dataset bias which is related to imbalanced data, ii) labeling bias that describes incorrect labels, and iii) contextual bias originating from image capture conditions.
Various bias mitigation strategies have been proposed to address these issues: a) pre-processing methods, which focus on data manipulation before training [23, 24], b) in-processing methods, which aim to reduce bias during the training phase [25, 26], and c) post-processing methods, which tend to mitigate bias after the model is trained [20, 21, 27]. Bias mitigation endeavors can target either the data or the algorithms themselves, depending on the source of bias.
@@
Despite the above-mentioned efforts to address bias, there is a notable research gap in automatic FER, particularly in in-the-wild datasets and novel deep models. In-the-wild datasets introduce biases from variations in illumination conditions, camera resolution, background noise, head pose, and gesture, as well as demographic attributes of age, gender, and race [28]. Moreover, many studies have focused on limited sources of bias, primarily examining label distributions or one or two demographic attributes [29, 30, 31]. This approach overlooks the multifaceted nature of bias in FER. Additionally, recent advancements in deep architectures, such as vision transformers [32] and large vision-language models [33, 34], have significantly impacted this field. However, their implications for bias and fairness in FER tasks remain largely unexplored. Fig. 1 shows the approach we utilize different metrics to evaluate various aspects of bias and fairness in the datasets and models.
@@
This research aims to address gaps by examining overlooked aspects of bias and fairness in FER. We provide a comprehensive analysis of bias and fairness in both datasets and deep models. Specifically, we argue that both datasets and models can introduce bias into FER systems. To assess the impact of datasets, we analyze four real-world FER datasets: AffectNet [35], ExpW [5], RAF-DB [36], and Fer2013 [37]. The bias in these datasets is evaluated using three metrics: Richness [38], Evenness [39], and Dominance [40]. Additionally, we investigate inherent biases in the datasets through experiments, as discussed in Sec. 3.1.
@@
In this study, we evaluate six deep models, commonly used for FER tasks. To this aim, we train three CNN models, i.e., MobileNetV1 [41], ResNet-50 [42], XceptionNet [43], and the vision-transformer-based ViT [32] model, using all the aforementioned datasets. From the two large language-vision models, CLIP [33] is fine-tuned using the datasets, while GPT-4o-mini [34] is evaluated without training or fine-tuning. We refer to MobileNetV1 as MobileNet and ResNet-50 as ResNet in the paper. To assess the bias of these models, we leverage four key bias metrics: Equalized Odds [44], Equal Opportunity [44], Demographic Parity [45, 46], and Treatment Equality [47]. These metrics capture aspects of bias and fairness, allowing for a comprehensive evaluation of the six models across multiple attributes. This methodology facilitates a fair and thorough comparison of their respective bias scores.
The experiments conducted in this study show that AffectNet is the most diverse dataset, while RAF-DB ranks the lowest in diversity. Regarding model performance, despite the high accuracy of the GPT-4o-mini and ViT models, they exhibited the highest bias scores when evaluated using the bias metrics. In contrast, ResNet demonstrated the lowest bias score, making it the fairest model among those tested in this study. Our implementation code and more detail about our experiments is reachable at GitHub.
@@ 2 Literature Review
FER has many applications, such as sensitive biometrics [48], where biased data or algorithms can cause irreparable damages. Understanding the sources of bias is essential for effectively addressing its challenges. Bias can originate from both datasets and models. Factors such as data collection and annotation processes, model architecture, objective functions, and evaluation methodologies are among the primary sources of bias. Figure 2 presents a general classification of bias sources. Addressing bias in deep machine learning methods can significantly enhance the reliability and accuracy of FER models. More importantly, developing fair and unbiased deep methods for FER methods potentially ensures equitable treatment across diverse demographic groups [49]. Thus, dealing with bias in FER has recently raised attention from researchers. Examining data distribution in existing FER datasets [50, 51, 52, 30, 53, 22], as well as analyzing the algorithms used in the FER models [28, 54, 21, 55, 27, 56] are the key approaches for studying bias in FER. In the following section, we review recent studies that address bias in FER models and datasets.
@@ 2.1 Bias in FER Datasets
Bias in datasets is broadly categorized into representational bias and stereotypical bias [50, 51, 52]. Representational bias focuses mainly on the demographic diversity of data, including age, gender, and race attributes. This bias is evaluated using metrics like Richness, Evenness, and Dominance (for definitions see Sec. 3.1). On the other hand, stereotypical bias investigates the correlation between specific subgroups and their corresponding labels, which can lead to unfair or biased outcomes.
Another categorization of bias in dataset divided them into intrinsic and extrinsic biases [57, 58, 59]. Intrinsic bias inherent to the process of data collection, i.e., the way data are collected, sampled, and labeled. Demographic attributes, such as age, gender, and race, as well as head-pose, gesture, eye-gaze, and expression intensity are studied as intrinsic biases. On the contrary, extrinsic bias refers to the disparities that could affect the model’s performance. For example, lab-controlled datasets can lead to high biases in models, because they are not collected based on the real-world scenarios. Extrinsic bias mainly stems from environmental conditions, such as camera resolution, background noise, illumination, and annotator bias. There are other definitions of bias in dataset. While Cheong et al. [22] grouped bias sources as observed and non-observed sources, Terhorst et al. [60] classified bias sources into demographic and non-demographic groups. Despite all these different bias categorizations, a strong consensus is that demographic attributes, including age, gender, and race, are the most effective sources of bias in dataset that need to be studied.
In the following we review the most notable works on this topic. Dominguez et al. [50] proposed seventeen representational metrics, where they ranked AffectNet [35], ExpW [5], RAF-DB [36], and Fer2013 [37] among the most diverse FER datasets. It is notable that their stereotypical metrics showed different results for this four datasets where they did not achieve good scores and rankings. Another research by Li et al. [30] studied the intrinsic bias among seven different FER datasets and reported that the main reasons for bias in datasets are cultural differences and data collection conditions. Huber et al. [53] studied bias in synthetic facial datasets where they revealed that synthetic datasets follow the same bias patterns as the authentic datasets used for training their corresponding models. Finally, Cheong et al. [22] used a directed acyclic graph (DAG) to explore the relation between sources of bias. They introduced gender and age as the strongest sources of bias in emotion labeling. These experiments highlight the importance of dataset analysis in FER tasks, especially demographic attributes and their correlation with the labels for each subgroup. In Sec. 3.1 we review some common metrics for data analysis, and in Sec. 4.1 demonstrate their results.
@@ 2.2 Bias in FER Algorithms
The analysis of bias in FER algorithms involves examining how deep models and algorithms impact various demographic groups, aiming to enhance fairness and improve emotion recognition accuracy. Below, we review recent studies addressing algorithmic bias. While some studies tried to reduce the bias of the model’s using data compensation, some experiments revealed that there are some inherent biases in the models irrelevant to the data.
To mitigate the effects of imbalanced data distribution, Yu et al. [54] utilized unlabeled data in a semi-supervised learning framework. This framework leveraged unlabeled facial data by assigning pseudo-labels to underrepresented samples in the datasets. As a result, their method achieved a more balanced data distribution and improved the overall accuracy of the FER system. In a related study, Kolahdouzi et al. [55] introduced a Hilbert-space kernel to compute the mean distribution of specific demographic groups. To reduce bias, they subtracted a facial image’s embedding from the extracted mean, obscuring sensitive attributes such as age, gender, and race. Furthermore, Liang et al. [27] investigated the influence of facial attributes on model performance. They categorized these attributes into protected attributes (e.g., gender and race) and unprotected attributes (e.g., pose and lighting). Their findings demonstrated that protected attributes affect model performance, contributing to bias in predictions.
In contrast, some researchers argue that biases originate from the models. Dooley et al. [21] hypothesized that bias stems from the architecture of deep models. They employed a joint CNN architecture with one component emphasizing fairness and another focusing on extracting hyperparameters to mitigate bias. Their results indicated that bias remains inherently tied to model architecture. Abdullah et al. [56] conducted experiments focusing on the attention areas allocated to faces by the models. They observed that the attention area assigned to dark-skinned faces was smaller compared to lighter skin tones, revealing an inherent bias in the models. Xu et al. [28] compared a baseline model, ResNet-18 [42], with two disentangled and attribute-aware models. Their experiments demonstrated that data augmentation did not significantly enhance the accuracy of the baseline model. Similarly, Amigo et al. [49] proposed a debiasing variational autoencoder to address bias in models. Their findings showed that, despite training on different datasets, models could perpetuate biases. Furthermore, Dominguez et al. [61] revealed that stereotypical biases are intrinsic to models, regardless of whether the training data is balanced or imbalanced. These findings underscore the importance of studying bias and fairness in both data and algorithms.
@@ 2.3 Perspectives
Bias and fairness in facial expression recognition are critical to preventing discrimination, ensuring equitable systems, improving performance across diverse populations, and upholding ethical standards. Real-world FER applications rely on in-the-wild datasets, making it imperative to thoroughly examine commonly used datasets to identify and address bias issues. While some studies have partially investigated these datasets, a more comprehensive exploration of popular FER datasets is needed.
At the same time, recent advancements in model architectures, such as deep convolutional neural networks, vision transformers, and large vision-language models, have significantly improved FER accuracy [62, 63, 64]. However, despite their broad adoption, these models must be rigorously evaluated for fairness and bias mitigation. This study provides an in-depth analysis of bias in four prominent FER datasets—AffectNet, ExpW, Fer2013, and RAF-DB—and assesses six state-of-the-art models, including MobileNet, ResNet, XceptionNet, ViT, CLIP, and GPT-4o-mini.
@@ 3.1 Data Analysis
In Sec. 2.1, we explored various classifications of bias, including intrinsic vs. extrinsic, observed vs. non-observed, demographic vs. non-demographic, and representational vs. stereotypical categories. Bias can manifest in many ways, such as through camera inconsistencies, background variations, body posture differences, and, most notably, demographic imbalances. Among these, demographic attributes—age, gender, and race—emerge as the most significant contributors to bias in FER datasets. As a result, these attributes, along with label distribution, warrant the most attention in this study to address bias and promote fairness.
Building on the taxonomy proposed by Dominguez et al. [50], this study analyzes bias in dataset using two major categories of metrics. Representational metrics examine the Richness, Evenness, and Dominance of the dataset, while stereotypical metrics reveal the local and global relationships between attributes.
Richness evaluates the total number of samples within a dataset. More samples lead to a higher Richness score. However, this metric solely considers the quantity of samples, without addressing their distribution across various attributes [38].
We consider $R_{i}=N_{i}$ as the basic Richness equation where $R_{i}$ is Richness of dataset $i$ and $N_{i}$ stands for the number of samples correspond to dataset $i$. To have a better understanding and scoring, we normalized the Richness score of each dataset, in the range $[0,1]$, through Eq. 1, where $max(N)$ is maximum Richness score of all the datasets.
$R_{i}=\frac{N_{i}}{max(N)},$ (1)
Evenness measures the equity of data distribution across different classes of attributes. It is commonly assessed using the Shannon’s Evenness Index [67], which quantifies how uniformly the samples are distributed in the datasets.
$E=-\frac{\sum_{i=1}^{C}p_{i}\log(p_{i})}{\log(C)},~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}p_{i}=\frac{n_{i}}{N}.$ (2)
Eq. 2 shows the Shannon’s Evenness Index, where $0\leq E\leq 1$ represents the Evenness, $C$ refers to the total number of classes, $n_{i}$ denotes the number of samples in class i, and $N$ is total number of samples. $E=1$ indicates a perfectly even distribution, whereas $E<1$ reflects an uneven data distribution.
Dominance measures the extent to which one group holds an advantage over other groups within a dataset. Since this definition gives more score to non-uniform data distribution, we subtracted it from 1, to convert it to the same range and the concept that we are using for the Richness and Evenness metrics. The proposed metric for assessing Dominance is the Simpson’s Dominance Index [40],
$D=1-\sum_{i=1}^{C}p_{i}^{2},~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}p_{i}=\frac{n_{i}}{N},$ (3)
which provides a quantitative evaluation of this phenomenon. In this equation, $0\leq D\leq 1$ represents Dominance, $C$ denotes the total number of classes, $p_{i}$ is the proportion of samples in each class $i$, $n_{i}$ and $N$ indicates the number of samples in a class, and the total number of samples, respectively. In this metric, a value of $D=0$ signifies that one class dominates entirely, while $D>0$ reflects a more evenly distribution across classes.
Global-Stereotypical and Local-Stereotypical metrics examine the correlation between an attribute and a class label. Global-Stereotypical metrics focus on the overall relationship between an attribute and data distribution, such as the connection between age groups and expression labels. Local-Stereotypical metrics delve into more specific relationships, analyzing the connection between a particular attribute value and a class label. For example, the relationship between the attribute value ’age-group = young’ and the Happy expression.
To investigate the local and global stereotypical metrics across different datasets, we extracted the data distribution for each dataset based on expression labels, as well as demographic attributes, age, gender, and race. The mathematical definition of the data distribution is provided in Eq. 4 as follows:
$P(A=a)=\frac{f_{a}}{N},$ (4)
where $A$ represents an attribute selected from expression, age, gender, or race, $a$ is any possible value for $A$, $f_{a}$ is the frequency of $a$, and $N$ is the total number of samples. In this study, utilizing DeepFace [68] definitions, age is classified into four ranges: [0$\sim$15], [16$\sim$32], [33$\sim$53], and [Over 54], gender is categorized as Woman or Man, and ultimately, six racial groups of White, Black, Asian, Indian, Latinx, and Middle Eastern, are considered.
In addition, we assessed the conditional data distribution of each dataset. Given that each expression class is associated with three attributes—age, gender, and race—we explored the data distribution under both single and joint conditions. This was done using the general distribution equation $P(A\mid B)=\frac{P(A\cap B)}{P(B)}$, where the denominator remains constant across all conditions, allowing us to simplify it to $P(A\mid B)=P(A\cap B)$. Equations 5,  6, and 7 present the mathematical formulation of this analysis.
@@
In Eqs. 567, $Y$ shows the expression, while $A$, $B$, $C$ correspond to the facial attributes age, gender, and race, with $y$, $a$, $b$, $c$ representing their respective values.

### v2

@@ Abstract
Building AI systems, including Facial Expression Recognition (FER), involves two critical aspects: data and model design. Both components significantly influence bias and fairness in FER tasks. Issues related to bias and fairness in FER datasets and models remain underexplored. This study investigates bias sources in FER datasets and models. Four common FER datasets—AffectNet, ExpW, Fer2013, and RAF-DB—are analyzed. The findings demonstrate that AffectNet and ExpW exhibit high generalizability despite data imbalances. Additionally, this research evaluates the bias and fairness of eight deep models, including three generic convolutional neural network (CNN) models: MobileNet, ResNet, XceptionNet, as well as three generic transformer-based models: ViT, CLIP, and GPT-4o-mini, plus two FER-built state-of-the-art models: POSTER and CEPrompt. Experimental results reveal that among the 6 generic models (exclude POSTER and CEPrompt), GPT-4o-mini and ViT achieve the highest accuracy scores, while they also display the highest levels of bias. On the other hand, considering all the models demonstrate that the best accuracies are achieved for two FER-specialized models (POSTER and CEPrompt), while they could not cope with the bias problem in FER task. These findings underscore the urgent need for developing new methodologies to mitigate bias and ensure fairness in datasets and models, particularly in affective computing applications. See our implementation details at https://github.com/MMHosseini/bias_in_FER.
@@
Bias, Fairness, Facial Expression Recognition, Facial Affect Analysis, Bias in FER Datasets, Bias in FER Models.
@@ 1 Introduction
Despite advancements in deep learning, bias and fairness issues in FER systems remain unexplored. In a general overview, bias refers to prejudices or favoritism toward groups, while fairness ensures impartiality and avoids disparities [9, 10]. Bias can be analyzed in two dimensions: datasets and models [9, 11, 12]. Dataset bias often arises from demographic imbalances, such as disparities in age, gender, and race [9]. Model bias stems from architecture, training process, and unrepresentative datasets [13, 14, 15, 16]. Bias impacts model performance for underrepresented groups, reduces generalizability, and compromises fairness. Therefore, studying bias in datasets and models is crucial for fairness in FER.
@@
There is a notable research gap in bias of automatic FER, particularly in in-the-wild datasets and novel deep models. In-the-wild datasets introduce biases from variations in illumination conditions, camera resolution, background noise, head pose, and gesture, as well as demographic attributes of age, gender, and race [17]. Many studies have focused on limited sources of bias in datasets, such as examining label distribution or one or two demographic attributes [18, 19, 20, 21], where they overlook the multifaceted nature of bias in FER. Additionally, recent advancements in deep architectures, such as vision transformers [22] and large vision-language models [23, 24], have significantly impacted this field. However, their implications for bias and fairness in FER tasks remain largely unexplored. These shortcomings prompted us to conduct an in-depth investigation into bias and fairness in facial expression recognition.
@@
This research aims to address gaps by examining overlooked aspects of bias and fairness in FER. We provide a comprehensive analysis of bias and fairness in both datasets and deep models. Specifically, we argue that both datasets and models can introduce bias into FER systems. Fig. 1 shows the mathematical metrics that we utilized to evaluate various aspects of bias and fairness in the datasets and models.
To assess the impact of datasets, we analyze four real-world FER datasets: AffectNet [25], ExpW [5], RAF-DB [26], and Fer2013 [27]. The bias in these datasets is evaluated using seven well-known statistical metrics that measure diversity, uncertainty, and dependence in the data. We also investigate bias through analyses of single and joint probability distributions over the labels and demographic attributes of age, gender, and race. In addition, we examine inherent bias by conducting two model-based experiments: (1) identifying which dataset each sample originates from, and (2) evaluating how well models trained on three datasets generalize to a held-out dataset.
@@
We study bias in models by evaluating eight popular deep models in FER. Specifically, we select three CNN models: MobileNetV1 [28], ResNet-50 [29], and XceptionNet [30]; three transformer-based models: ViT [22], CLIP [23], and GPT-4o-mini [24]; and two state-of-the-art FER-specialized models: POSTER [31] and CEPrompt [32]. We assess bias in these models using four key metrics that quantify how predictions or errors vary across demographic groups, capturing disparities in outcomes, opportunities, and relative treatment and evaluating overall fairness. By measuring multiple aspects of bias, these metrics provide a comprehensive evaluation across different attributes, allowing a fair and thorough comparison of the models’ bias scores. In summary, our main contributions on this research are:
  • Comprehensive bias assessment framework: a unified pipeline that analyzes both dataset and model bias in FER, with cross-dataset generalization tests and detailed demographic evaluations.
  • Advanced dataset-bias metrics: introduction of seven statistical and information-theoretic measures together with micro- and macro-level correlation analyses to reveal distributional and conditional dependencies.
  • Extensive model-bias benchmarking: fairness evaluation of eight representative CNN, Transformer, and FER-specialized models using four established fairness metrics, providing a reproducible benchmark for future studies.
Our implementation code and more detail about our experiments is reachable at GitHub.
@@ 2 Literature Review
Bias in machine learning can be classified in various ways. Pagano et al. [33] introduced three types of bias: pre-existing bias, technical bias, and emergent bias. Mehrabi et al. [9] provided a broader taxonomy including historical, representation, measurement, aggregation, and evaluation biases. Understanding the sources of bias is essential for effectively addressing its challenges. Bias can originate from both datasets and models. Factors such as data collection, cultural interpretation, annotation process, model architecture, objective functions, and evaluation methodologies are among the primary sources of bias. Developing fair and unbiased deep methods for FER potentially ensures equitable treatment across diverse demographic groups [34]. Thus, dealing with bias in FER has recently raised attention from researchers. In the following, we review recent studies on bias in FER datasets and models.
@@ 2.1 Bias in FER Datasets
Bias in datasets refers to systematic skews or unfairness in the data that can distort model training and predictions [35]. Several taxonomies have been proposed to categorize dataset bias. One common distinction is between representational bias and stereotypical bias[21, 36, 37]. The former targets demographic diversity, while the latter examines correlations between subgroups and labels. Another categorization separates intrinsic and extrinsic biases[38, 39, 40]: intrinsic bias arises from the data collection process and demographic or behavioral factors, whereas extrinsic bias stems from external conditions such as camera resolution, illumination, background noise, or annotator subjectivity. Additional dimensions have also been identified, with Cheong et al. [41] distinguishing among data collection bias, labeling bias, and contextual bias, and Jones et al. [42] categorizing bias into prevalence disparities, presentation disparities, and annotation disparities. Despite these diverse perspectives, there is strong consensus that demographic attributes, particularly age, gender, and race, remain the most influential sources of dataset bias and warrant the most attention.
Although establishing a taxonomy of dataset bias is important, understanding the underlying sources of bias provides guidance for mitigation. Based on our study, we categorize sources of bias into five main classes: 1) demographic disparities, referring to imbalances in age, gender, and racial groups that can affect model fairness across different populations; 2) environmental conditions, capturing inconsistencies in illumination, background scenes, and camera quality that may influence visual perception; 3) positional variations, including differences in body gestures, eye-gaze, head-pose, and clothing, which can alter the appearance of expressions; 4) expression diversity, reflecting imbalances between labels, and also easily recognizable and more subtle or challenging expressions that can bias model learning; and 5) annotation orientation, addressing cultural and subjective differences in labeling, which may introduce inconsistencies in ground truth data. A visual summary of these bias sources is provided in Fig. 2.
In the matter of in-the-wild FER datasets analysis, Dominguez et al. [21] found that while they are diverse in terms of demographic representation, they still exhibit weaknesses under stereotypical bias metrics. Li et al. [19] analyzed intrinsic bias across seven FER datasets, identifying cultural differences and data collection conditions as major contributors. Huber et al. [43] showed that synthetic facial datasets replicate the bias patterns present in the authentic datasets used for training. Cheong et al. [41] employed directed acyclic graphs to examine bias sources, highlighting gender and age as the most influential factors in emotion labeling. Collectively, these studies emphasize the critical role of dataset analysis in FER tasks, particularly the impact of demographic attributes on label distributions. Subsequent sections review common metrics for data analysis (Sec. 3.1) and present experimental results (Sec. 4.1).
@@ 2.2 Bias in FER Algorithms
Model bias is a systematic error in decision-making that produces unfair outcomes [12]. Another view describes it as biased results caused by inadequate model specifications [44]. Several taxonomies of bias in ML models have been proposed. Mehrabi et al. [9] identify algorithmic bias, evaluation bias, and optimization bias as key forms: algorithmic bias reflects the model’s assumptions, evaluation bias stems from benchmarks favoring certain groups, and optimization bias arises when objectives or loss functions benefit majority groups. Binns et al. [45] instead classify bias as direct, indirect, and structural. Direct bias occurs when algorithms deliberately treat groups unfairly to boost accuracy, indirect bias emerges from accidental errors when models fail to capture features, and structural bias results from broader organizational inequalities shaping algorithmic behavior.
Similar to dataset bias, we identify five main sources of bias in ML models: (1) algorithmic bias, which occurs when the model overfits to a dominant group in the training data; (2) model design bias, arising from choices in architecture, objective functions, loss functions, or training strategies; (3) feature extraction bias, when the model places undue emphasis on certain facial attributes; (4) generalization bias, where the model performs well only on a specific dataset but fails to transfer to others; and (5) evaluation bias, which stems from benchmarks or metrics that overlook fairness considerations (see Fig. 2).
Recent studies of bias in ML algorithms reveal that unfair outcomes in FER often persist even when data issues are addressed, suggesting that models themselves introduce bias. Dooley et al. [46] showed that bias can stem directly from network architecture, while Abdullah et al. [47] found that attention mechanisms allocate less focus to darker-skinned faces. Xu et al. [17] reported that data augmentation offers little improvement when the model is inherently biased, and Amigo et al. [34] demonstrated that models can reproduce bias across datasets despite debiasing strategies. Dominguez et al. [48] further argued that stereotypical biases are intrinsic to models, regardless of data balance. Together, these findings highlight that algorithmic bias is deeply embedded in model architectures and learning mechanisms, not just in the data.
@@ 2.3 Perspectives
Bias and fairness in facial expression recognition (FER) are essential for preventing discrimination, ensuring equity, improving performance across diverse groups, and maintaining ethical standards. Because real-world FER relies on in-the-wild datasets, it is crucial to examine these datasets to identify and address sources of bias. Although some studies have investigated them, a more systematic exploration with undiscovered metrics will open new insights into the inherent bias in the datasets. At the same time, rapid advancements in model architectures have significantly boosted FER accuracy [49, 50, 51]. Yet, despite these improvements, questions of fairness and bias in such models remain underexplored.
In this work, we conduct a comprehensive analysis of four widely used FER datasets (AffectNet, ExpW, Fer2013, and RAF-DB) using evaluation metrics. In parallel, we assess eight state-of-the-art models, including six general-purpose architectures (MobileNet, ResNet, XceptionNet, ViT, CLIP, and GPT-4o-mini) and two FER-specialized models (POSTER and CEPrompt), to investigate how bias evolves with model advancements. Taken together, this is the first systematic effort to provide an in-depth evaluation of both dataset- and model-level biases in FER.
@@ 3 Methodology
TABLE I: Taxonomy of the equations used in this research.
Symbol Meaning Example
$A,B,C$ Symbol of each attribute AA: Age, $B$: Gender
$a,b,c$ Specific demographic group $a$: Man, $b$: Woman
$Y$ Expression set (labels) -
$y$ Specific expression value $y$: Happy
$H$ Entropy $H(Y)$: Entropy of labels
$n$ Number of labels (7 in our study) -
$k$ Number of demographic groups 4 age groups
METMET Models’ bias metrics -
ATTATT Attributes (age, gender, race) -
$p$ Probability -
@@ 3.1 Data Analysis
Bias in data manifest in ways, such as camera inconsistencies, background variations, body posture differences, and, most notably, demographic imbalances. Among these, demographic attributes, i.e., age, gender, and race, emerge as the most significant contributors to bias in FER datasets. As a result, these attributes, along with label distribution, warrant the most attention in this study to address bias and promote fairness.
To evaluate demographic bias in our datasets, we employ Wasserstein Distance (WD), Jensen–Shannon Divergence (JSD), Conditional-Entropy Bias Index (CEBI), Sparsity Index (SI), Normalized Shannon Entropy (NSE), Normalized Label Skewness (NLS), and Group Normalized Mutual Information (GNMI). Unlike the metrics in [21], these measures directly capture distributional differences and conditional dependencies that influence machine-learning models. WD and JSD quantify divergence in label distributions across demographic groups, revealing subtle shifts that broad diversity scores may miss. CEBI and GNMI measure information gain about labels from demographic attributes, making them sensitive to conditional or interaction-based bias. NSE and NLS offer normalized, interpretable indicators of entropy and skewness, enabling clear cross-dataset comparisons. We customized all the metrics in a way to work as a bias score in the range of $[0,1]$, where 0 means no bias and 1 shows the maximum bias.
Because our metrics draw on probability and information theory, they relate more directly to model training and predictive fairness than the ecological variety measures in [21]. This connection clarifies how demographic imbalance affects classifier performance and guides mitigation strategies. While classical diversity metrics help describe overall variety, the selected measures provide a complementary, machine-learning-focused view of dataset bias. Next, we present details for each metric. Table I presents the taxonomy of our equations.
Wasserstein Distance (WD), also known as Earth Mover’s Distance, is a metric sensitive to imbalance magnitude and class structure. For example, in a similar distribution of data for two attributes Male and Female, WD is equal to 0, while for two skewed and dissimilar distribution it gives a high value. This metric captures how much ”probability mass” should be shifted to reach to an equal data distribution:
@@
Jensen–Shannon Divergence (JSD) quantifies the similarity between two probability distributions. While Wasserstein Distance quantifies bias by measuring the minimal effort required to transport one group’s distribution to another, Jensen–Shannon Divergence captures differences in their probabilistic information content without accounting for geometric distance. JSD is equal to 0 when two distributions are similar, and 1 when they are disjoint. In FER it determines how different the labels are distributed across demographic groups:
$\begin{aligned} JSD=\frac{2}{k(k-1)}\sum_{a,b\in A}\frac{1}{2n}\sum_{i=1}^{n}\left(H(a,m)+H(b,m)\right),\\ H(a,m)=p(y_{i}|A=a)\log\frac{p(y_{i}|A=a)}{m(y_{i})},\quad\quad\quad\\ m(y_{i})=\frac{p(y_{i}|A=a)+p(y_{i}|A=b)}{2}.\quad\quad\quad\quad\end{aligned}$
Conditional-Entropy Bias Index (CEBI) measures how an emotion class depends on a specific attribute in a group. For example, how the Male gender influences the Happy expression. A low CEBI means that the attribute strongly determines the label. To provide a consistent scale in which 0 indicates no bias and 1 indicates maximum bias, we normalized its equation and then reversed by subtracting it from 1:
$\begin{aligned} CEBI=\frac{1}{k}\sum_{a\in A}1-\frac{H(Y\mid a)}{H(Y)}.\end{aligned}$
Sparsity Index (SI) is a normalized version of Gini coefficient and quantifies the imbalance data distribution across the dataset in the range of [0, 1]. Unlike WD and JSD, which assess differences between group distributions, Sparsity Index measures how concentrated or uneven the label distribution is within each group, capturing intra-group imbalance rather than inter-group distance. For example, it evaluates how even the expressions are distributed on Male gender, independent of the Female gender. A higher SI reveals that only a few demographic groups dominate the samples, making it a clear indicator of representational bias:
$\begin{aligned} SI=\frac{1}{k}\sum_{a\in A}\frac{\left|\|\mathbf{p}\|_{2}^{2}-\tfrac{1}{n}\right|}{1-\tfrac{1}{n}},\quad\quad\quad\|\mathbf{p}\|_{2}^{2}=\sum_{y\in Y}p(y)^{2}.\end{aligned}$
Normalized Shannon Entropy (NSE) summarizes the entire demographic distribution into a single value without requiring pairwise group comparisons. NSE measures the overall label diversity within a group, providing a scale-invariant view of internal balance. For instance, it can evaluate how fairly the Happy expression is distributed across various racial groups; if it is biased toward an overrepresented group, NSE assigns a higher bias score. This metric simultaneously captures both the richness and evenness of group representation:
$\begin{aligned} NSE=\frac{1}{k}\sum_{a\in A}p(A=a)\left|1-\frac{H(Y\mid a)}{\log n}\right|.\end{aligned}$
Normalized Label Skewness (NLS) quantifies the asymmetry of class label distributions within each demographic group. For example, for the Male group, it measures how much each expression’s distribution deviates from the overall average. While Normalized Shannon Entropy (NSE) captures overall label diversity and balance, NLS highlights the directional imbalance of over- or underrepresented classes within a group. Large positive or negative values indicate strong bias, so NLS is normalized to the range [0, 1] for comparability:
$\begin{aligned} NLS=\frac{1}{k}\sum_{a\in A}\frac{|S|}{1+|S|},\quad\quad S=\frac{1}{n}\sum_{i=1}^{n}\left[\frac{p(y_{i}|A=a)-\mu}{\sigma}\right]^{3}.\end{aligned}$
Group-Normalized Mutual Information (GNMI) measures non-linear statistical dependencies between labels and demographic groups. It quantifies how much information about a label can be inferred from group membership. For example, it evaluates the relationship between the Black racial group and the Happy expression, capturing how much knowing the demographic group reduces uncertainty about the label. While CEBI measures conditional label uncertainty given the group, GNMI normalizes this dependency for scale-independent comparison. High GNMI values indicate that the label is highly predictable from the demographic attribute:
$\begin{aligned} GNMI(Y,A)=\frac{I(Y;A)}{\sqrt{H(Y)H(A)}},\quad\quad\quad\\ I(Y;A)=\sum_{a\in A}\sum_{y\in Y}p(y,a)\log\frac{p(y,a)}{p(y)p(a)},\quad\end{aligned}$
Macro-Level Correlation Bias and Micro-Level Correlation Bias metrics examine the correlation between an attribute and a class label. Macro-Level Correlation Bias focuses on the overall relationship between an attribute and data distribution, such as the connection between ageage attribute and expressionexpression labels. On the other hand, Micro-Level Correlation Bias delves into analyzing the connection between a particular group and a class label. For example, the relationship between the group $age\mathord{=}[0\mathord{\sim}15]$ and $expression\mathord{=}Happy$.
To investigate Macro-Level Correlation Bias, we extracted the probability distributions of expressions and attributes in each dataset, denoted by $p(Y\mathord{=}y)$ and $p(A\mathord{=}a)$, respectively. To study Micro-Level Correlation Bias, given that each expression class is associated with three attributes—age, gender, and race—we explored the data distribution under single, double, and triplet conditions, denoted by $p(Y\mathord{=}y|A\mathord{=}a)$, $p(Y\mathord{=}y|A\mathord{=}a,B\mathord{=}b)$, and $p(Y\mathord{=}y|A\mathord{=}a,B\mathord{=}b,C\mathord{=}c)$, respectively.